Solve for the unknown interest rate in each of the following: (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
  

Present Value Years Interest Rate Future Value
$ 160 4 % $ 192
280 18 733
31,000 19 128,034
30,261 25 350,964
Answer #1

FV = $192 = $160(1 + r)4

r = ($192 / $160)1/4 – 1 = 4.66%

FV = $733 = $280(1 + r)18

r = ($733 / $280)1/18 – 1 = 5.49%

FV = $128,034 = $31,000(1 + r)19

r = ($128,034 / $31,000)1/19 – 1 = 7.75%

FV = $350,964 = $30,261(1 + r)25

r = ($350,964 / $30,261)1/25 – 1 = 10.30%